OpenStudy (anonymous):

find the x-intercept, y-intercept and the vertical asymptote of g(x)=ln(x+1)

myininaya (myininaya):

x-intercept set y=0 (g(x)=y)) 0=ln|x+1| 1=x+1 x=0 so the x intercept is (0,0) so the y-intercept is (0,0) also

OpenStudy (dumbcow):

vertical asymptote: x+1 = 0 x = -1

myininaya (myininaya):

right cow :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

x-intercept, y=0 0=ln(x+1) 1=x+1 x=0, y-intercept, x=0 g(x)=ln(1) =0 vertical asymptote, when x is not in the domain, since \[x \le-1\] is not in the domain, so the asymptote is when x=-1